\begin{thebibliography}{AGKR24}

\bibitem[Aar05]{aaronson2005quantum}
Scott Aaronson.
\newblock Quantum computing, postselection, and probabilistic polynomial-time.
\newblock {\em Proceedings of the Royal Society A: Mathematical, Physical and
  Engineering Sciences}, 461(2063):3473--3482, 2005.
\newblock \href {https://doi.org/10.1098/rspa.2005.1546}
  {\path{doi:10.1098/rspa.2005.1546}}.

\bibitem[Aar06]{aaronson_subtle}
Scott Aaronson.
\newblock Oracles are subtle but not malicious.
\newblock In {\em Proceedings of the 21st Annual IEEE Conference on
  Computational Complexity}, pages 340–--354. IEEE Computer Society, 2006.
\newblock \href {https://doi.org/10.1109/CCC.2006.32}
  {\path{doi:10.1109/CCC.2006.32}}.

\bibitem[Aar07]{aaronson2007learnability}
Scott Aaronson.
\newblock The learnability of quantum states.
\newblock {\em Proc. R. Soc. A.}, 463(2088):3089--3114, 2007.
\newblock \href {https://doi.org/10.1098/rspa.2007.0113}
  {\path{doi:10.1098/rspa.2007.0113}}.

\bibitem[Aar17]{aaronsonBlogErrors2017}
Scott Aaronson.
\newblock Yet more errors in papers, May 2017.
\newblock Accessed 14 Jan. 2024.
\newblock URL: \url{https://scottaaronson.blog/?p=3256}.

\bibitem[AB09]{arora2009computational}
Sanjeev Arora and Boaz Barak.
\newblock {\em Computational Complexity: a Modern Approach}.
\newblock Cambridge University Press, 2009.

\bibitem[AD14]{aaronson2014full}
Scott Aaronson and Andrew Drucker.
\newblock A full characterization of quantum advice.
\newblock {\em SIAM Journal on Computing}, 43(3):1131--1183, 2014.
\newblock \href {https://doi.org/10.1137/110856939}
  {\path{doi:10.1137/110856939}}.

\bibitem[AGKR24]{agarwal2024quantum}
Avantika Agarwal, Sevag Gharibian, Venkata Koppula, and Dorian Rudolph.
\newblock Quantum polynomial hierarchies: {K}arp-{L}ipton, error reduction, and
  lower bounds, 2024.
\newblock \href {https://arxiv.org/abs/2401.01633} {\path{arXiv:2401.01633}}.

\bibitem[Fen03]{fenner2003pp}
Stephen~A. Fenner.
\newblock {PP}-lowness and a simple definition of {AWPP}.
\newblock {\em Theory of Computing Systems}, 36:199--212, 2003.
\newblock \href {https://doi.org/10.1007/s00224-002-1089-8}
  {\path{doi:10.1007/s00224-002-1089-8}}.

\bibitem[FR99]{fortnow1999complexity}
Lance Fortnow and John Rogers.
\newblock Complexity limitations on quantum computation.
\newblock {\em Journal of Computer and System Sciences}, 59(2):240--252, 1999.
\newblock \href {https://doi.org/10.1006/jcss.1999.1651}
  {\path{doi:10.1006/jcss.1999.1651}}.

\bibitem[FSW09]{fortnow2009fixed}
Lance Fortnow, Rahul Santhanam, and Ryan Williams.
\newblock Fixed-polynomial size circuit bounds.
\newblock In {\em Proceedings of the 24th Annual IEEE Conference on
  Computational Complexity}, pages 19--26. IEEE, 2009.
\newblock \href {https://doi.org/10.1109/CCC.2009.21}
  {\path{doi:10.1109/CCC.2009.21}}.

\bibitem[GLV24]{GLV24oblivious}
Karthik Gajulapalli, Zeyong Li, and Ilya Volkovich.
\newblock Oblivious classes revisited: Lower bounds and hierarchies.
\newblock ECCC: TR24-049, 2024.
\newblock URL: \url{https://eccc.weizmann.ac.il/report/2024/049/}.

\bibitem[GM15]{goldreich2015input}
Oded Goldreich and Or~Meir.
\newblock Input-oblivious proof systems and a uniform complexity perspective on
  {P}/poly.
\newblock {\em ACM Transactions on Computation Theory}, 7(4):1--13, 2015.
\newblock \href {https://doi.org/10.1145/2799645} {\path{doi:10.1145/2799645}}.

\bibitem[KL80]{karp1980some}
Richard~M. Karp and Richard~J. Lipton.
\newblock Some connections between nonuniform and uniform complexity classes.
\newblock In {\em Proceedings of the 12th annual ACM Symposium on Theory of
  Computing}, pages 302--309, 1980.
\newblock \href {https://doi.org/10.1145/800141.804678}
  {\path{doi:10.1145/800141.804678}}.

\bibitem[Kup15]{kuperberg2015hard}
Greg Kuperberg.
\newblock How hard is it to approximate the {J}ones polynomial?
\newblock {\em Theory of Computing}, 11(1):183--219, 2015.

\bibitem[LFKN92]{lund1992algebraic}
Carsten Lund, Lance Fortnow, Howard Karloff, and Noam Nisan.
\newblock Algebraic methods for interactive proof systems.
\newblock {\em Journal of the ACM (JACM)}, 39(4):859--868, 1992.
\newblock \href {https://doi.org/10.1145/146585.146605}
  {\path{doi:10.1145/146585.146605}}.

\bibitem[Li93]{li1993counting}
Lide Li.
\newblock {\em On the counting functions}.
\newblock PhD thesis, The University of Chicago, 1993.
\newblock URL:
  \url{https://www.proquest.com/dissertations-theses/on-counting-functions/docview/304080357/se-2}.

\bibitem[LP17]{levin2017markov}
David~A. Levin and Yuval Peres.
\newblock {\em {M}arkov chains and mixing times}, volume 107.
\newblock American Mathematical Soc., 2017.

\bibitem[MN16]{morimae16}
Tomoyuk Morimae and Harumichi Nishimura.
\newblock Quantum interpretations of {AWPP} and {APP}.
\newblock {\em Quantum Info. Comput.}, 16(5–-6):498–--514, 2016.
\newblock \href {https://doi.org/10.26421/QIC16.5-6-6}
  {\path{doi:10.26421/QIC16.5-6-6}}.

\bibitem[MVW99]{miltersen1999super}
Peter~Bro Miltersen, N.~V. Vinodchandran, and Osamu Watanabe.
\newblock Super-polynomial versus half-exponential circuit size in the
  {E}xponential {H}ierarchy.
\newblock In {\em International Computing and Combinatorics Conference}, pages
  210--220. Springer, 1999.
\newblock \href {https://doi.org/10.1007/3-540-48686-0_21}
  {\path{doi:10.1007/3-540-48686-0_21}}.

\bibitem[MW05]{marriott2005quantum}
Chris Marriott and John Watrous.
\newblock Quantum {A}rthur--{M}erlin games.
\newblock {\em Computational Complexity}, 14:122--152, 2005.
\newblock \href {https://doi.org/10.1007/s00037-005-0194-x}
  {\path{doi:10.1007/s00037-005-0194-x}}.

\bibitem[Vin05]{vinodchandran2005note}
N.~V. Vinodchandran.
\newblock A note on the circuit complexity of {PP}.
\newblock {\em Theoretical Computer Science}, 347(1):415--418, 2005.
\newblock \href {https://doi.org/10.1016/j.tcs.2005.07.032}
  {\path{doi:10.1016/j.tcs.2005.07.032}}.

\bibitem[Wat08]{watrous2008quantum}
John Watrous.
\newblock Quantum computational complexity, 2008.
\newblock \href {https://arxiv.org/abs/0804.3401v1} {\path{arXiv:0804.3401v1}}.

\bibitem[Wil14]{williams2014nonuniform}
Ryan Williams.
\newblock Nonuniform {ACC} circuit lower bounds.
\newblock {\em Journal of the ACM (JACM)}, 61(1):1--32, 2014.
\newblock \href {https://doi.org/10.1145/2559903} {\path{doi:10.1145/2559903}}.

\bibitem[Zoo]{ZooBQPmpoly}
Complexity {Z}oo: {BQP/poly}, {BQP/mpoly}, {BQP/qpoly}, {BQP}.
\newblock Accessed 13 Mar. 2024.
\newblock URL: \url{https://complexityzoo.net/Complexity_Zoo:B}.

\end{thebibliography}

\documentclass[11pt,letterpaper,onecolumn]{article}

\pdfoutput=1

\usepackage[letterpaper, lmargin=1.25in, rmargin=1.25in, tmargin=1in, bmargin=1in]{geometry}
\usepackage[utf8]{inputenc}
\usepackage[english]{babel}

\usepackage{amsmath,amssymb,amsthm,amsfonts,latexsym,bbm,xspace,thm-restate}
\usepackage{graphicx,float,mathtools,braket,mathdots}
\usepackage{enumitem,booktabs,forest,soul,comment}
\usepackage[useregional]{datetime2}
\DTMusemodule{english}{en-US}
\usepackage[colorlinks,citecolor=blue,,linkcolor=magenta,bookmarks=true]{hyperref}
\usepackage[nameinlink,capitalize]{cleveref}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{corollary}[theorem]{Corollary}

\crefname{fact}{Fact}{Facts}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}

\let\Pr\relax
\DeclareMathOperator{\Pr}{Pr}
\DeclareMathOperator{\E}{E}

\DeclareMathOperator{\poly}{poly}

\newcommand{\nats}{\mathbb N}

\newcommand{\class}[1]{\ensuremath{\mathsf{#1}}\xspace}
\mathchardef\mhyphen="2D %

\newcommand{\PTIME}{\class{P}}
\newcommand{\BPP}{\class{BPP}}
\newcommand{\NP}{\class{NP}}

\newcommand{\AC}{\class{AC}}

\newcommand{\NEXP}{\class{NEXP}}

\newcommand{\PSPACE}{\class{PSPACE}}

\newcommand{\BQP}{\class{BQP}}
\newcommand{\QCMA}{\class{QCMA}}
\newcommand{\QMA}{\class{QMA}}

\newcommand{\Pii}[1][i]{\class{\Pi^{p}_{#1}}}

\newcommand{\GapP}{\class{GapP}}
\newcommand{\sharpP}{\class{\#P}}
\newcommand{\PP}{\class{PP}}

\newcommand{\CH}{\class{CH}}
\newcommand{\CHi}[1][i]{\class{C_{#1}P}}

\newcommand{\Ppoly}{\class{P\!/\!poly}}
\newcommand{\BQPpoly}{\class{BQP\!/\!poly}}
\newcommand{\BQPqpoly}{\class{BQP\!/\!qpoly}}
\newcommand{\YQP}{\class{YQP}}
\newcommand{\YQPpoly}{\class{\YQP\!/\!poly}}
\newcommand{\YQPstar}{\class{YQP^*}}
\newcommand{\YQPstarpoly}{\class{\YQP^*\!/\!poly}}
\newcommand{\AWPP}{\class{AWPP}}
\newcommand{\APP}{\class{APP}}

\newcommand{\postBQP}{\class{postBQP}}

\newcommand{\prob}[1]{\textsc{#1}}

\newcommand{\calB}{\mathcal{B}}

\newcommand{\bbI}{\mathbb{I}}

\newcommand{\abs}[1]{\left\lvert #1 \right\rvert}

\newcommand{\ie}{i.e.\xspace}

\newcommand{\eg}{e.g.\xspace}

\newcommand{\badv}{b_{\text{adv}}}
\newcommand{\bout}{b_{\text{out}}}

\title{Even quantum advice is unlikely to solve \texorpdfstring{$\PP$}{PP}}

\author{Justin Yirka\thanks{Supported via Scott Aaronson by a Simons Investigator Award.}
	\\\small{\textsl{The University of Texas at Austin}}\\\small{\texttt{\href{mailto:yirka@utexas.edu}{yirka@utexas.edu}}}
}

\date{May 2024}

\begin{document}

\maketitle

\begin{abstract}
	We give a corrected proof that
	if $\PP\subseteq \BQPqpoly$, then the Counting Hierarchy collapses,
	as originally claimed by [Aaronson 2006].
	This recovers the related unconditional claim that $\PP$ does not have circuits
	of any fixed size $n^k$ even with quantum advice.
	We do so by proving that $\YQPstar$, an oblivious version of $\QMA\cap\class{coQMA}$,
	is contained in $\APP$, and so is $\PP$-low.
\end{abstract}

\section{Introduction}

Do reasonably-sized circuits solve hard problems, given that we allow the computation
to vary (non-uniformly) as the problem size grows?
While directly answering this question for classes such as $\NP$ has proven difficult,
progress has been made showing conditional results,
such as the Karp-Lipton theorem that if $\NP\subseteq \Ppoly$,
then the polynomial hierarchy collapses \cite{karp1980some},
or showing upper bounds against larger classes,
such as $\PP$ or $\NEXP$ \cite{vinodchandran2005note,williams2014nonuniform}.
Exploring further, we can consider quantum computation.
In this model, circuits are typically uniformly generated
but might accept non-uniform \emph{advice} strings as part of their input.
Moreover, quantum circuits can not only receive classical
advice strings ($\BQPpoly$), but
quantum \emph{advice states} ($\BQPqpoly$).

In \cite{aaronson_subtle}, Aaronson proved new quantum circuit lower bounds,
among other results.
In particular, he gave several results characterizing quantum circuits' ability
to solve problems in the class $\PP$,
the class of problems decidable by probabilistic algorithms with no promise gap,
\ie{} which accept with probability at least $1/2$ or strictly less than $1/2$.
$\PP$ contains both $\BPP$ and $\NP$ and is contained in $\PSPACE$.
Aaronson proved that $\PTIME^\PP$ does not have circuits of size $n^k$ for any
fixed constant $k$ even if the circuits use quantum advice states.
Second, he claimed a quantum analogue of the Karp-Lipton theorem, showing that if
$\PP\subseteq \BQPqpoly$, then the Counting Hierarchy ($\CH$) collapses to $\QMA$,
where the Counting Hierarchy is the infinite sequence of classes $\CHi[1]=\PP$ and
$\CHi[i]=\left(\CHi[i-1]\right)^\PP$,
and where $\QMA$ is a quantum analogue of $\NP$.
Similarly, he showed that under the stronger assumption $\PP \subseteq \BQPpoly$,
using classical advice instead of quantum advice,
then $\CH = \QCMA$.
Third, Aaronson combined these results to give the unconditional bound that $\PP$
does not have classical or quantum circuits of size $n^k$ for
any fixed constant $k$ even with quantum advice.\footnote{Slightly earlier,
Vinochandran \cite{vinodchandran2005note} gave a proof that $\PP$ does not
have \emph{classical} circuits of fixed polynomial size.}

However, Aaronson later noted there was an error in one of the proofs \cite{aaronsonBlogErrors2017}.
The first of the above results was unaffected,
but the proof of the second result only held under the stronger assumption that
$\PP\subseteq \BQPpoly$.
This also meant the third result only held for quantum circuits with classical,
not quantum, advice.
Fortunately, no other results in \cite{aaronson_subtle} were affected,
but no fix for this bug
was forthcoming.

Very briefly, the error was a claim that for oracle classes of the form
$\class{C}^{\BQPqpoly}$, if a machine for the base class $\class{C}$
is able to find the quantum advice state that will be used by the oracle machine,
then the base machine can ``hard-code'' the advice state into its oracle queries so that
the oracle no longer needs the power to find its own advice,
thus reducing $\class{C}^{\BQPqpoly}$ to $\class{C}^{\BQP}$.
This approach works for classes with classical advice, like $\class{C}^{\BQPpoly}$.
But, because complexity classes and their associated oracles
are defined in terms of (classical) strings as input,
there is no way to hard-code a general quantum advice state into a query.

In this note, we give a corrected proof of Aaronson's full claims.
We show that if $\PP\subseteq \BQPqpoly$,
then the Counting Hierarchy collapses to $\QMA$ and in fact to $\YQPstar$.
Given this correction, Aaronson's proof for the third claim,
that $\PP$ does not have circuits of size $n^k$
for any fixed constant $k$ even with quantum advice, now goes through.

Our primary technical contribution is to show $\YQPstar \subseteq \APP$.
Here, $\YQPstar$ is an oblivious version of $\QMA\cap\class{coQMA}$, meaning
there exists a useful proof state which depends only on the size of the input.
Crucially, a $\YQPstar$ protocol includes a proof-verification circuit that tests
if the given quantum state is a ``good'' proof, before the proof is used
to determine whether a particular input is a YES or NO instance.
We apply the in-place error reduction technique of
Marriott and Watrous \cite{marriott2005quantum}
to this proof-verification circuit.
The class $\APP$ is a subclass of $\PP$ with an arbitrarily small but nonzero promise gap.
It is known to have the nice property that $\PP^\APP = \PP$ \cite{li1993counting}.
Thus, $\YQPstar$ is also $\PP$-low.
Our corrected proof combines this result with
the known equality $\BQPqpoly = \YQPstarpoly$,
serendipitously proven by Aaronson with Drucker \cite{aaronson2014full}.
Now, instead of following Aaronson's original attempt to collapse $\PP^\PP$ to $\PP^\BQPqpoly$ to $\PP^\BQP$ to $\PP$,
we can collapse $\PP^\PP$ to $\PP^\YQPstarpoly$ to $\PP^\YQPstar$ to $\PP$.

\sloppy %
Our results provide stronger implications and improved bounds for quantum circuits
with quantum advice
and establish new insights into
$\PP$-lowness and classes within $\APP$.
Compared to other quantum Karp-Lipton style bounds, including that if
$\QCMA\subseteq \BQPpoly$, then $\class{QCPH}$ collapses \cite{agarwal2024quantum}
and that if
$\NP\subseteq \BQPqpoly$, then $\Pii[2] \subseteq \QMA^{\class{PromiseQMA}}$ \cite{aaronson2014full},
the supposition $\PP\subseteq \BQPqpoly$
and the implied collapse of $\CH$ are both formally stronger.
As for unconditional bounds, following Aaronson's unaffected result that
$\PTIME^\PP$ does not have quantum circuits with quantum advice of any fixed
polynomial size, our corrected result bound against $\PP$ is the first improved bound
of fixed-size circuits with quantum advice.
Regarding $\PP$-lowness, our primary lemma establishes $\YQPstar$ as the largest
natural quantum complexity class known to be $\PP$-low,
improving on the fact that $\BQP$ is $\PP$-low \cite{fortnow1999complexity}.\footnote{Morimae and Nishimura \cite{morimae16}
	gave definitions involving quantum postselection constructed to equal $\AWPP$
	and $\APP$, which are $\PP$-low.}
Finally, for $\APP$, while the largest witness-based class previously known
to be contained in $\APP$ was $\class{FewP}$ \cite{li1993counting},
our result shows that $\APP$ in fact contains oblivious-witness classes
including $\YQPstar\supseteq\class{YMA^*}\supseteq\class{YP^*}\supseteq\class{FewP}$.

\section{Preliminaries}

In this section, we discuss non-uniform circuits,
give definitions for $\YQP$ and $\APP$,
discuss $\PP$ and $\GapP$,
and finally state a fact relating quantum circuits to $\GapP$.
For a deeper introduction to these classes and other concepts,
see \cite{aaronson_subtle,aaronson2014full} and \eg{} \cite{arora2009computational}.

\paragraph{\texorpdfstring{~}{Non-uniform circuits}}
The classes of non-uniform circuits we consider, including $\Ppoly, \BQPpoly, \BQPqpoly$,
share the following key characteristics.
First, they are defined in terms of circuits or advice that may depend
on the size of the problem input (but not on the input itself),
with no requirement that the circuit or advice is generated by a uniform algorithm.
Second, the circuits are defined with bounded fan-in and fan-out, in contrast to
classes such as $\AC_0$ or $\class{QAC}_0$.
Third, the classes consider circuits of polynomial-size,
where the size is the number of gates in a circuit.

Advice is typically considered ``trusted'' in that there is no promised behavior
when given the ``wrong'' advice,
so an analysis of correctness can usually assume the ``right'' advice is provided.
In line with this convention,
we assume the standard definitions of $\BQPpoly$ and $\BQPqpoly$
in which a circuit is only required to accept with high or low probability
(outside of the ``promise gap'') when the correct advice is provided,
although the same notation has sometimes been used to refer to other definitions,
see \eg{} \cite{ZooBQPmpoly}.

\paragraph{\texorpdfstring{~}{YQP}}
The class $\YQP$ was first described in \cite{aaronson2007learnability},
but the definition was later corrected by Aaronson and Drucker \cite{aaronson2014full}.
Informally, it is the oblivious version of $\QMA{}\cap\class{coQMA}$,
so that the witness sent by Merlin depends only on the length of the input.
In contrast to the advice of $\Ppoly$, this has been described as
``untrusted advice'' \cite{aaronson2007learnability}.
Oblivious proofs can also be thought of as restricting non-uniform classes,
like $\Ppoly$ or $\BQPqpoly$,
to advice which is verifiable~\cite{goldreich2015input}.

\begin{definition}\label{def:YQP}
	A language $L$ is in $\YQP$ if there exists a polynomial-time uniform family of
	quantum circuits $\{Y_n\}_{n\in \nats}$ that satisfy the following.
	Circuit $Y_n$ is of size $\poly(n)$ and takes as input
	$x\in \{0,1\}^n$, a $p(n)$-qubit state $\rho$ for some $p(n)\leq \poly(n)$,
	and an ancilla register initialized to the all-zero state,
	and has two designated 1-qubit ``advice-testing'' and ``output'' qubits.
	$Y_n(x,\rho)$ acts as follows:
	\begin{enumerate}
		\item First, $Y_n$ applies a subcircuit $A_n$ to all registers, after which
		the advice-testing qubit is measured, producing a value $\badv\in\{0,1\}$.
		\item Next, $Y_n$ applies a second subcircuit $B_n$ to all registers,
		then measures the output qubit, producing a value $\bout\in\{0,1\}$.
	\end{enumerate}
	These output bits satisfy the following:
	\begin{itemize}
		\item For all $n$, there exists a $\rho_n$ such that for all $x$,
		the advice bit satisfies $\E[\badv] \geq 9/10$.
		\item For any $x,\rho$ such that $\E[\badv]\geq 1/10$,
		on input $x,\rho$ we have
		\[
			\Pr\left[\bout = L(x) \mid \badv = 1\right] \geq 9/10 .
		\]
	\end{itemize}
	$L$ is in the subclass $\YQPstar$ if the family can be chosen such that $\badv$ is
	independent of $x$.
\end{definition}

Just as $\class{Oblivious\mhyphen{}NP}$ is unlikely to contain $\NP$ \cite{fortnow2009fixed},
it also seems unlikely that $\QMA$ is contained in $\YQP$.
On the other hand, it is straightforward to show that any sparse language can be
verified obliviously,
so $\class{FewP}\subseteq \class{YP^*}$ and $\class{FewQMA}\subseteq \YQPstar$.
We also have the trivial bounds $\BQP\subseteq \YQPstar \subseteq \YQP\subseteq \QMA$
and $\YQP\subseteq \BQPqpoly$.
Studying $\YQP$ may be motivated by the use of oblivious complexity classes
in constructing circuit lower bounds \cite{fortnow2009fixed,GLV24oblivious},
by the fact that $\BQPqpoly = \YQPstarpoly = \YQPpoly$ shown by \cite{aaronson2014full},
or by the results shown in this work.

\paragraph{\texorpdfstring{~}{APP}}
The class $\APP$ was introduced by Li \cite{li1993counting} in pursuit of
a large class of $\PP$-low languages.
We use the equivalent definition given by Fenner \cite[Corollary 3.7]{fenner2003pp}.

\begin{definition}\label{def:APP}
	$L\in \APP$ if and only if there exist
	functions $f,g\in \GapP$ and constants $0\leq \lambda < \upsilon \leq 1$ such that
	for all $n$ and $x\in\{0,1\}^n$,
	we have $g(1^n)>0$ and
	\begin{itemize}
		\item If $x\in L$ then
		$\upsilon g(1^n) \leq f(x) \leq g(1^n)$;
		\item If $x\notin L$ then $0\leq f(x) \leq \lambda g(1^n)$.
	\end{itemize}
\end{definition}

In the above definition, recall that $\GapP$ is the closure of $\sharpP$ under
subtraction.
In other words, while every function $f\in \sharpP$ corresponds to a nondeterministic
polynomial-time Turing Machine $N$ such that $f(x)$ equals the number of accepting paths
of $N(x)$,
a $\GapP$ function equals the number of accepting paths minus the number
of rejecting paths.

$\APP$ is a subclass of $\PP$ and is $\PP$-low, meaning $\PP^\APP = \PP$.
Recall that $\PP$ can be thought of as comparing a $\sharpP$ function
to a threshold exactly, with no promise gap.
The class in fact remains unchanged if it is defined as comparing a $\GapP$ function
to a threshold, and the threshold may be as simple as one-half of the possible paths or as complex as a
$\GapP$ function.
In these terms,
$\APP$ can be thought of as comparing a $\GapP$ function (here $f(x)$) to some threshold
(here $g(1^n)$),
where the complexity of the threshold is limited to a $\GapP$ function which may
depend on the input size but not the input,
and where there is some arbitrarily small but nonzero promise gap
(from $\lambda g(1^n)$ to $\upsilon g(1^n)$).

The best known upper bound on $\APP$ is $\PP$.
Compared with the class $\class{A_0PP}=\class{SBQP} \subseteq \PP$ \cite{kuperberg2015hard},
$\class{A_0PP}$ contains $\QMA$ and is \emph{not} known to be $\PP$-low,
while $\APP$ is not known to contain even $\NP$ but is $\PP$-low.

\paragraph{\texorpdfstring{~}{Useful fact}}
We will use the following fact shown for uniform circuit families by Watrous
\cite[Section~IV.5]{watrous2008quantum}, and shown earlier for QTMs by Fortnow and
Rogers \cite{fortnow1999complexity}.

\begin{lemma}\label{lem:quantumGapP}
	For any polynomial-time uniformly generated family of quantum circuits
	$\{Q_n\}_{n\in \nats}$ each of
	size bounded by a polynomial $t(n)$,
	there is a $\GapP$ function $f$ such that for all $n$-bit $x$,
	\[
		\Pr\left[Q_n(x)\text{ accepts}\right] = \frac{f(x)}{5^{t(n)}} .
	\]
\end{lemma}

\section{Results}

We first prove our main technical result, that $\YQPstar\subseteq \APP$,
improving on the largest proof-based complexity class known to be contained
in $\APP$.

Our approach is as follows.
$\APP$ evaluates the ratio of two $\GapP$ functions, where one of the functions
is only allowed to depend on the input length. By \cref{lem:quantumGapP}, functions
in $\GapP$ can encode the output probabilities of quantum circuits.
So, for a $\YQPstar$ computation with circuit $Y$ and subcircuit $A$,
we run them both on the maximally-mixed state
and ask $\APP$ to determine the ratio of their acceptance probabilities.
In other words, when $A$ accepts, however large or small that probability is,
does $Y$ usually accept or usually reject?
We require the error reduction technique of
Marriott and Watrous \cite{marriott2005quantum} to make the error in $A$ negligible.

\begin{lemma}\label{lem:YQPinAPP}
	$\YQPstar\subseteq \APP$.
\end{lemma}
\begin{proof}
	Consider any language $L\in \YQPstar$.
	Let $\{Y_n,A_n,B_n\}_{n\in\nats}$ be the associated family of circuits
	and subcircuits,
	in which $Y_n$ takes string $x$ and a supposed witness or advice state as input,
	in which subcircuit
	$A_n$ validates the advice and produces output bit $\badv$,
	and in which,
	given $A_n$ accepted, $B_n$ uses the advice to verify whether the particular input
	$x$ is in $L$, producing the output bit $\bout$.
	Note that because we consider $\YQPstar$, the circuit $A_n$ only takes
	the witness state, not $x$, as input.
	Let $k$ and $m$ be polynomials in $n$ denoting the respective sizes of the ancilla
	and proof registers.

	We use the technique of strong, or in-place, error reduction of Marriott and Watrous
	\cite{marriott2005quantum} on the circuits $A_n$
	with a polynomial $q$ in $n$ of our choosing
	to produce a new circuit family $\{A'\}_{n\in\nats}$
	such that for any proof $\rho$,
	\begin{itemize}
		\item $\Pr\left[A_n(\rho)\right] \geq \frac{9}{10} \Rightarrow
		\Pr\left[A'_n(\rho)\right]\geq 1-2^{-q}$;
		\item $\Pr\left[A_n(\rho)\right] \leq \frac{1}{10} \Rightarrow
		\Pr\left[A'_n(\rho)\right]\leq 2^{-q}$.
	\end{itemize}
	For later use, we choose $q > \max\{3m,10\}$.

	Recall the error reduction algorithm of \cite{marriott2005quantum} involves,
	given some quantum input or witness state,
	applying a circuit $C$,
	recording whether the output is $\ket{0}$ or $\ket{1}$ in a variable $y_i$,
	applying $C^{\dagger}$, recording whether the circuit's ancilla register is
	in the all-zero
	state or not in a variable $y_{i+1}$, and repeating these steps for some number of
	iterations $M$.
	Call the full, amplified circuit $C'$.

	Applying the error-reduction procedure,
	we define $\{A''_n\}_{n\in\nats}$ to be the amplified
	circuits $\{A'_n\}_{n\in\nats}$ with the additional rule that the circuit
	accepts iff both $\badv=1$ and the final two recorded variables $y_{2M}=y_{2M+1}=1$.
	Further, define $\{A'''_n\}_{n\in\nats}$ so that $A'''_n = A''_n(\frac{\bbI}{2^m})$,
	with the maximally mixed state hard-wired into the proof register.
	Similarly, we define $\{Y'_n\}_{n\in\nats}$ to apply the amplified subcircuit
	$A'_n$ and $B_n$,
	we define $\{Y''_n\}_{n\in\nats}$ to apply $A''_n$ and $B_n$ and thus accept iff
	$\badv,\bout,y_{2M},y_{2M+1}$ all equal~1,
	and we define $\{Y'''_n\}_{n\in\nats}$ so that $Y'''_n(x) = Y''_n(x,\frac{\bbI}{2^m})$
	with the maximally mixed state hard-wired into the proof register,
	meaning that it uses $A'''_n$ as a subcircuit.

	\paragraph{Error-reduction properties}
	We remark on what the bits $y_{i}$ tell us about the state of the circuit.
		Studying the proof of \cite{marriott2005quantum},
		if after applying $C^\dagger$, a recorded bit $y_{2i+1}=1$,
		then the state of the ancilla register is projected into the all-zero state.
		Now, suppose the circuit $C'$ is applied to an $m$-qubit proof state,
		so there are $2^m$ eigenstates $\{\ket{\lambda_i}\}_{i\in [2^n]}$ of $C'$.
		Further studying the proof of \cite{marriott2005quantum},
		if the initial state given to $C'$ was an eigenstate $\ket{\lambda_i}$,
		and after a round of applying $C^\dagger$ the recorded bit is $y_{2i+1}$,
		then not only is the ancilla register known to be in the all-zero state,
		but the final state of the proof register is the same as its initial state,
		$\ket{\lambda_i}$.

	We are also able to characterize the probabilities of these outcomes.
		Intuitively, consecutive bits are transitions which depend on whether we expect
		the circuit beginning with a properly initialized all-zero ancilla register to
		produce an output qubit close to $\ket{1}$ or to $\ket{0}$, and vice-versa.
		Suppose an eigenstate $\ket{\lambda_i}$ is accepted by the
		original circuit $C$ with probability $p$.
		Then when $C'$ is run on $\ket{\lambda_i}$,
		we have that $\Pr\left[y_{2i+1}=1 \mid y_{2i}=1\right] = p$.
		Next, we consider $\Pr\left[y_{2i}=1\right]$.
		We may analyze this probability with a two-state Markov chain,
		such that if $y_{2i}=1$, then $y_{2i+2}=1$ with probability $p^2 + (1-p)^2$
		and $y_{2i+2}=0$ with probability $2p(1-p)$,
		while if $y_{2i}=0$, then $y_{2i+2}=0$ with probability $p^2+(1-p)^2$ and
		$y_{2i+2}=1$ with probability $2p(1-p)$.
		If we suppose $p > 1/2$,
		and given $C'$ begins in the all-zero state ($y_0=1$),
		then it is straightforward to conclude that for any particular $i$,
		we have $\Pr\left[y_{2i}=1\right]> 1/2$
		(see \eg{} \cite[Example 1.1]{levin2017markov}).

	\paragraph{Analysis}
	Applying \cref{lem:quantumGapP}, there exist $\GapP$ functions $f,g$ and polynomials
	$r,t$ such that for all $n$-bit $x$,
	\begin{gather*}
		\Pr\left[A'''_n\text{ accepts}\right] = \frac{f(1^n)}{5^{r(n)}}
		\quad\text{and}\quad
		\Pr\left[Y'''_n(x)\text{ accepts}\right] = \frac{g(x)}{5^{t(n)}} .
	\end{gather*}
	The function $f$ depends only on the input length $n$, not $x$,
	because the circuit $A'''_n$ is independent of $x$.
	Next, we define $F(1^n) = f(1^n) 5^{t(n)-r(n)}$, which is a $\GapP$ function
	since $5^{t(n)-r(n)}\in \class{FP}\subseteq \GapP$ and $\GapP$ is closed under
	multiplication.
	Given the definition of $\YQPstar$ guarantees there
	exists a ``good'' proof for circuit $A_n$,
	we have  $f(1^n),F(1^n)>0$.
	Combining these definitions,
	\begin{gather*}
		\frac{g(x)}{F(1^n)} =
		\frac{\Pr\left[Y'''_n(x)\text{ accepts}\right]}
		{\Pr\left[A'''_n\text{ accepts}\right] } .
	\end{gather*}

	We will show bounds on the ratio $g(x)/F(1^n)$ based on whether $x$ is in $L$ or
	not in $L$ in order to prove $L$ is in $\APP$.
	First, note that the ratio is upper-bounded by 1 since $Y'''_n$ only accepts if the
	subcircuit $A'''_n$ accepts,
	and it is lower-bounded by 0 since probabilities are non-negative.
	Next, let $\left\{\ket{\lambda_i}\right\}_{i\in[2^m]}$ be the set of
	eigenvectors $\ket{\lambda_i}$ of the circuit $A_n$.
	By writing the maximally mixed state, which is hard-wired into the proof register of
	$Y'''_n$,
	in terms of this eigenbasis, we find
	\begin{align*}
		& \frac{\Pr\left[Y'''_n(x)\text{ accepts}\right]}
		{\Pr\left[A'''_n\text{ accepts}\right]}
		= \frac{\Pr\left[Y''_n(x,\frac{\bbI}{2^m})\text{ accepts}\right]}
		{\Pr\left[A''_n(\frac{\bbI}{2^m})\text{ accepts}\right]}
		= \frac{\sum\nolimits_{i=1}^{2^{m}}
			\Pr\left[Y''_n(x,\ket{\lambda_i})\text{ accepts}\right]}
		{\sum\nolimits_{i=1}^{2^{m}} \Pr\left[A''_n(\ket{\lambda_i})
			\text{ accepts}\right]} \\
		=& \frac{\sum\nolimits_{i=1}^{2^{m}}
			\Pr\left[Y''_n(x,\ket{\lambda_i})\text{ accepts}
				\mid A''_n(\ket{\lambda_i})\text{ accepts} \right]
			\Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right]}
		   {\sum\nolimits_{i=1}^{2^{m}} \Pr\left[A''_n(\ket{\lambda_i})
		   	\text{ accepts}\right]}    \\
		=& \frac{\sum\nolimits_{i=1}^{2^{m}}
			\Pr\left[B_n(x,\ket{\lambda_i})\text{ accepts}\right]
			\cdot \Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right]}
		  {\sum\nolimits_{i=1}^{2^{m}} \Pr\left[A''_n(\ket{\lambda_i})
		  	\text{ accepts}\right]}
	\end{align*}
	where we have used the fact that $Y''_n$ accepting requires that
	$A''_n$ accepts and our observation
	that $A''_n$ accepting guarantees the initial eigenstate $\ket{\lambda_i}$ is sent
	on to the subcircuit $B_n$ within $Y''_n$.
	Define
	\[
		\calB = \left\{ i\in[2^m] \mid \Pr\left[A_n(\ket{\lambda_i})\right]
		\leq 0.1 \right\} ,
	\]
	which are intuitively the ``bad'' proofs, such that states in $\calB$ will be rejected
	by $A'_n$ with high probability while the ``not bad'' states in $\overline{\calB}$
	cause $B_n$ to output the correct answer with high probability.
	Then we can rewrite both the numerator and denominator in the above ratio to give
	\[
		\frac{
			\text{\small\ensuremath{\sum\nolimits_{i\in \calB} }}
			\text{\footnotesize\ensuremath{\Pr\left[B_n(x,\ket{\lambda_i})\text{ accepts}\right]
				\cdot \Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right] + }}
			\text{\small\ensuremath{\sum\nolimits_{i\in\overline{\calB}} }}
			\text{\footnotesize\ensuremath{ \Pr\left[B_n(x,\ket{\lambda_i})\text{ accepts}\right]
				\cdot \Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right] }}
		}
		{
			\sum\nolimits_{i\in \calB} \Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right]
			+ \sum\nolimits_{i\in\overline{\calB}} \Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right]
		} .
	\]
	We will use this expression as the starting point for our analysis of the
	YES and NO cases.

	Now, suppose we have a YES instance with $x\in L$.
	We are guaranteed at least one proof is accepted by $A$ with
	high probability, and denote it by $\ket{\lambda^*}$.
	Then, we may calculate that $g(x)/F(1^n)$ is at least
	\begin{align*}
		\frac{
			\sum\nolimits_{i\in \calB} 0
			+ \sum\nolimits_{i\in\overline{\calB}} \frac{9}{10}
			\Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right]
		}
		{
			\sum\nolimits_{i\in \calB} 2^{-q}
			+ \sum\nolimits_{i\in\overline{\calB}}
			\Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right]
		}
		&= \frac{
			\sum\nolimits_{i\in\overline{\calB}} \frac{9}{10}
			\Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right]
		}
		{
			\abs{\calB} 2^{-q}
			+ \sum\nolimits_{i\in\overline{\calB}} \Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right]
		} \\
		&\geq \frac{\frac{9}{10} \Pr\left[A''_n(\ket{\lambda^*})\text{ accepts}\right]}
		{ \abs{\calB} 2^{-q} + \Pr\left[A''_n(\ket{\lambda^*})\text{ accepts}\right]} ,
	\end{align*}
	where the second line follows by the fact that $x/(c+x)$ decreases as $x$ decreases.
	Next, we use the same fact, the earlier bound on the probability
	that the error-reduction variables $y_{2M},y_{2M+1}$ equal 1,
	and our choice $q> \max\{3m,10\}$
	to find that the above is at least
	\begin{align*}
	  	\frac{\frac{9}{10}(1-2^{-q}) (0.9)(0.5) }
	  	{\abs{\calB}2^{-q} + (1-2^{-q}) (0.9)(0.5)}
		&\geq \frac{ 0.405 (1-2^{-q}) }{2^{m-q} + 0.45(1-2^{-q}) } \\
		&\geq \frac{0.405(1-2^{-q}) }{2^{-q/3} + 0.45(1-2^{-q}) } \\
		&\geq \frac{ 0.405 (1 - 2^{-10})}{ 2^{-10/3} + 0.45 (1-2^{-10})}
		> 0.73 .
	\end{align*}

	On the other hand, consider a NO instance.
	We have that $g(x)/F(1^n)$ is at most
	\begin{align*}
		\frac{
			\abs{\calB} 2^{-q} + \sum_{i\in\overline{\calB}}
			\frac{1}{10} \Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right]
		}
		{
			\sum_{i\in \calB} 0 + \sum_{i\in \overline{\calB}} \Pr\left[A''_n(\ket{\lambda_i})\text{ accepts}\right]
		}
		&\leq
		\frac{2^{m-q}}
		{\sum_{i\in \overline{\calB}} \Pr\left[A''_n(\ket{\lambda_i})
			\text{ accepts}\right]
		}
		+
		\frac{1}{10} \\
		&\leq
		\frac{2^{m-q}}{2^{-m}(1-2^{-q})} + \frac{1}{10}	\\
		&
		< \frac{2^{-q/3} }{1-2^{-q}} + \frac{1}{10}
		< 0.2	,
	\end{align*}
	where the second line follows by the existence of $\ket{\lambda^*}$
	and the final line uses our choice of $q > \max\{3m,10\}$.

	We have shown a constant separation of $g(x)/F(1^n)$ in YES and NO instances.
	This satisfies the definition of $\APP$ in \cref{def:APP} of $\APP$,
	so we conclude $\YQPstar\subseteq \APP$.
\end{proof}

Next, the fact $\APP$ is known to be $\PP$-low \cite[Theorem 6.4.14]{li1993counting}
gives us the following corollary.

\begin{corollary}\label{lem:YQPisLow}
	$\YQPstar$ is $\PP$-low, \ie $\PP^{\YQPstar}=\PP$.
\end{corollary}

For intuition, an alternative proof of \cref{lem:YQPisLow} might have relied
on the equality $\PP=\postBQP$ \cite{aaronson2005quantum},
where $\postBQP$ has the ability to post-select, \ie{} it is guaranteed to output
the correct answer with high probability \emph{conditioned on} some other event
which may occur with very small probability.
So, instead of $\PP^\YQPstar$, we might have considered $\postBQP^\YQPstar$.
Whenever the $\postBQP$ machine would make a query, it instead could
run the $\YQPstar$ proof-validation circuit on
the maximally mixed state, post-select on it accepting,
then simulate the rest of the $\YQPstar$ computation.

We are now able to give a corrected proof of the result originally claimed for
$\BQPqpoly$ but only proved for $\BQPpoly$ by Aaronson \cite{aaronson_subtle}.
We mostly repeat Aaronson's proof, but substitute $\YQPstar$ where he relied on $\QMA$.

\begin{theorem}\label{thm:main}
	If $\PP \subseteq \BQPqpoly$, then the Counting Hierarchy collapses to
	$\CH = \QMA = \YQPstar$.
\end{theorem}

\begin{proof}
	Suppose $\PP \subseteq \BQPqpoly$. From \cite{aaronson2014full}, we know that
	$\BQPqpoly=\YQPstarpoly$.
	Then in $\YQPstar$, without any trusted advice, Arthur can request Merlin
	sends many copies of the quantum advice $\ket{\psi}$ and a description of the
	circuit $C$ such that $C,\ket{\psi}$ compute \prob{Permanent},
	a $\PP$-complete problem.
	Of course, this advice is now untrusted.
	Arthur verifies that $C,\ket{\psi}$ in fact work on a large fraction of inputs
	by simulating the interactive protocol for $\sharpP$ due to \cite{lund1992algebraic},
	which also works for $\PP$, using $C,\ket{\psi}$ in place of the prover.
	If the protocol accepts (meaning the ``prover'' worked),
	then Arthur can use the random self-reducibility of
	\prob{Permanent} to generate a circuit $C'$ which is correct on \emph{all} inputs
	(see \eg{} \cite[Sec. 8.6.2]{arora2009computational}).
	Thus, we have $\PP = \YQPstar$.

	In this way, any level of the Counting Hierarchy
	$\CHi[i] = \left(\CHi[i-1]\right)^\PP$ with $i>1$ is reducible to
	$\left(\CHi[i-1]\right)^{\YQPstar}$
	which by \cref{lem:YQPisLow} equals $\CHi[i-1]$.
	This works recursively for all levels, collapsing $\CHi[i]$ to $\CHi[1]=\PP$,
	so that all of $\CH=\PP=\YQPstar$.
\end{proof}

Given the above result, we can also fully recover the following result originally
claimed by Aaronson \cite{aaronson_subtle}, giving an improved unconditional upper bound
on fixed-size quantum circuits with quantum advice.

\begin{theorem}\label{thm:PPcircs}
	$\PP$ does not have quantum circuits of size $n^k$ for any fixed $k$.
	Furthermore, this holds even if the circuits can use quantum advice.
\end{theorem}
\begin{proof}
	Suppose $\PP$ does have circuits of size $n^k$.
	This implies $\PP\subseteq \BQPqpoly$, which by \cref{thm:main} implies
	$\CH = \YQPstar$, which includes
	$\PTIME^\PP = \PP = \YQPstar$.
	Together, there are circuits of size $n^k$ for $\PTIME^\PP$, which contradicts
	the result of \cite[Theorem 4]{aaronson_subtle} (unaffected by the bug)
	that $\PTIME^\PP$ does not have such circuits even with quantum advice.
\end{proof}

In fact, \cite{aaronson_subtle} noted
that the proof showing
$\PTIME^\PP$ does not have circuits of size $n^k$ for fixed $k$
even with quantum advice can be strengthened.
Substituting this stronger result into the above proof,
we have that \cref{thm:PPcircs} can be strengthened to show for all functions
$f(n)\leq 2^n$,
the class $\class{PTIME(f(f(n)))}$,
which is like $\PP$ but for machines of running time $f(f(n))$,
requires quantum circuits using quantum advice of size at least $f(n)/n^2$.
In particular, this implies $\class{PEXP}$, the exponential-time version of $\PP$,
requires quantum circuits with quantum advice of ``half-exponential'' size (meaning
a function that becomes exponential when composed with itself \cite{miltersen1999super}).

\bibliographystyle{alphaurl}
\bibliography{bibliography.bib}

\end{document}